24. The equation 12x² + 4kx + 3 = 0 has real and equal roots, if
(a) k = ±3
(b) k = ±9
(c) k = 4
(d) k = ±2

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: a Explaination:Reason: Here a = 12, b = 4k, c = 3 Since the given equation has real and equal roots ∴ b² – 4ac = 0 ⇒ (4k)² – 4 × 12 × 3 = 0 ⇒ 16k² – 144 = 0 ⇒ k² = 9 ⇒ k = ±3